provide a proof of spontaneity of the following reaction
4Fe(s)+ 3O2(g)->2Fe2O3(s)

provide a proof of spontaneity of the following re
Answer #1

4Fe(s) + 3O2(g) -------> 2Fe2O3(s)

delta G0rkn = 2*(delta G0f Fe2O3) - 4*(delta G0f Fe) - 3*(delta G0f O2) = -2*(-741.0) - 4*(0) - 3*(0) = -1482 kJ

Since, delta G0rkn < 0 ; therefore the reaction is spontaneous.